how to i work this problem out?
if my weight is 201 lb and i want to loss 5% of that, what and how do i get the answer?

Found 2 solutions by stanbon, unlockmath:
Answer by stanbon(75887)

how to i work this problem out?
if my weight is 201 lb and i want to loss 5% of that, what and how do i get the answer?
----
201-(0.05)(201) = 0.95*201 = 190.95 lbs
---
You want to lose 0.05*201 = 10.05 lbs
=========================================
Cheers,
Stan H.

Answer by unlockmath(1688)

Hello,
Here's one way to do it.
Multiply .05 times 201 and we get:
10.05 pounds
Make sense?
RJ
www.math-unlock.com